Prepare Job-Order Cost Sheets, Predetermined Overhead Rate, Ending Balance of WIP, Finished Goods, and COGS

At the beginning of June, Donewell Company had two jobs in process, Job 44 and Job 45, with the following accumulated cost information:

Job 44 Job 45
Direct materials $5,000 $1,800
Direct labour 4,400 3,520
Applied overhead 3,800 3,040
Balance, June 1 $13,200 $8,360

During June, two more jobs (46 and 47) were started. The following direct materials and direct labour costs were added to the four jobs during the month of June:

Job 44 Job 45 Job 46 Job 47
Direct materials $2,500 $7,110 $1,800 $1,700
Direct labour 800 6,400 900 560

At the end of June, Jobs 44, 45, and 47 were completed. Only Job 45 was sold. On June 1, the balance in Finished Goods was zero.

Required:

1. Calculate the overhead rate based on direct labour cost. Round the percentage to the nearest whole number. For example, .6372 would be rounded to 64%.
%

2. Prepare a brief job-order cost sheet for the four jobs. Show the balance as of June 1 as well as direct materials and direct labour added in June. Apply overhead to the four jobs for the month of June, and show the ending balances. Round your answers to the nearest whole dollar. For those boxes in which no entry is required, enter "0".

Donewell Company
Job-Order Cost Sheet
For the Month of June
Job 44 Job 45 Job 46 Job 47
Beginning balance, June 1 $ $ $ $
Direct materials
Direct labour
Applied overhead
Total, June 30 $ $ $ $

3. Calculate the ending balances of Work in Process and Finished Goods as of June 30. Round your answers to the nearest dollar.

Work in process $
Finished goods $

4. Calculate the Cost of Goods Sold for June. Round your answer to the nearest dollar.
$
Answer #1

1) overhead rate = 3800/4400 = 86%

2) Calculate following

Donewell Company
Job-Order Cost Sheet
For the Month of June
Job 44 Job 45 Job 46 Job 47
Beginning balance, June 1 $13200 $8360 $ $
Direct materials 2500 7110 1800 1700
Direct labour 800 6400 900 560
Applied overhead 688 5504 774 482
Total, June 30 $17188 $27374 $3474 $2742

3. Calculate the ending balances of Work in Process and Finished Goods as of June 30. Round your answers to the nearest dollar.

Work in process $3474
Finished goods 17188+2742 = 19930

4) Cost of goods sold = 27374